# Sales Whisperer: A Human-Inconspicuous Attack on LLM Brand Recommendations

### v1

@@ 4.1 Observations on Paraphrased Prompts
To promote a target brand, we apply this intuition. We explore attacking LLMs by testing paraphrased prompts. Specifically, we generate a set of prompts asking for recommendations of products in a certain category that are all paraphrases of one another. Our approach to this is described in §5.2. We show that these various paraphrased prompts (§6.1), although similar, can lead to increased prominence of a brand in the LLM’s responses. Given enough paraphrases, we’re able to find a prompt that causes LLMs to recommend a target brand more often.
This approach does not require any access to LLMs’ internal weights or token probabilities. However, it has a caveat, we need to try many paraphrases and collect many responses to find an optimal prompt. This computational cost might be an economic burden for the adversaries. Additionally, when we generate the paraphrases, some are likely less similar to each other and thus human-detectable. While we are able to find paraphrases that result in a large change in the probability that a certain brand is mentioned, even while using only a few prompts per category, as we describe in §5.2, in cases where an adversary does not have the time or computational resources to collect responses to multiple paraphrased prompts, adversaries may want to take another approach. As such, we propose another method to perturb prompts. To reduce computational overhead, with this new approach we assume access to the logits of LLMs in §4.2.
4.2 Synonym-Replaced Adversarial Prompts
The fundamental idea of this attack is to generate prompts by replacing words with synonyms and picking one that minimizes a loss function, reducing the need to collect computationally expensive responses. As we described in §2.2, while many NLP domain attacks have been proposed to be inconspicuous to humans, they either do not have accompanying user studies or are shown to be conspicuous to humans, motivating our approach. With this attack, we propose an improved synonym replacement method to generate human-inconspicuous perturbations.
Existing synonym replacement methods [23] either use WordNet, an established synonym dictionary, or a self-defined synonym dictionary [25, 26]. Nouns and verbs were found to have various forms (e.g., plural forms of nouns and verb tenses) [27], and thus an adjective synonym dictionary was suggested [28]. However, we found that synonyms of adjectives listed in existing dictionaries might not be synonyms in the context of brand recommendations. For example, WordNet and the existing adjective synonym dictionary both suggested “raw” as a synonym for “newest,” while it is human-conspicuous to recommend “the raw smartphone” instead of “the newest smartphone.” Thus, we created a new synonym dictionary compatible with brand recommendations. Our synonym dictionary mainly consists of adjectives but also includes other parts of speech while keeping the tense of verbs consistent between synonyms. In our dictionary, “select” and “choose” are synonyms, creating a synonym group, and other tenses of these verbs are excluded. “Exact,” “accurate,” and “precise” also creats a synonym group, this time of three. Our dictionary contains a total of 94 words in 36 synonym groups. Each word may have at most seven synonyms.
@@ 4.2 Synonym-Replaced Adversarial Prompts
Since paraphrasing prompts imposes an economic burden (§4.1), systematically testing and gathering responses for all candidate perturbed prompts generated by synonym replacement would have a high cost. One prompt we tested had six words with synonyms and therefore candidate perturbed prompts with the same meaning. Rather than testing all candidates, we used a loss-guided candidate-selection method. Prior work suggested that when adversaries have white-box access to LLMs (i.e., when adversaries can access internal architecture and weights of LLMs), adversaries may use gradient search to perturb prompts [7]. Compared to prior work, however, our method does not require access to model weights or gradients. Our search space is much smaller, we do not need consider all possible tokens a model accepts, just synonyms [7]. Instead, we computed a logit-based loss for all possible combinations of synonym replacements and picked the combination with the lowest loss. It should be noted, that other candidates than the one with the lowest could be used as well, and we used this method to narrow the candidates down in order to be able to test more base prompt and target brand combinations.
@@
Specifically, we use the following loss function. Using the same notation as existing works [7], we considered LLMs as a mapping from a sequence of tokens to a distribution over the next token . In other words, LLMs generate a probability . The probability that LLMs generate a sequence can be denoted as . In contrast to Zou et al., that aim to generate a specific sequence of tokens using the loss function
@@
Intuitively, our loss function aims to cause LLMs to generate some sequence from the set right after the prompt . As we describe in §3, adversaries may not need to cause LLMs to mention a specific brand name to recommend the brand. For example, causing LLMs to recommend either “Macbook” or “Apple” meets adversaries’ goal to recommend the brand “Apple” for the category “laptops”, and in this case, might be .
@@
Our loss function aims to increase the likelihood of words to appear from the target set right after the prompt, and we use the increased likelihood of words from this target set as a proxy to estimate how close adversaries are to their goal of promoting a brand (described in §3). We use increased likelihood to estimate which combination of a prompt’s synonym replacement may cause LLMs to recommend the target brand more often. Notably, despite the loss-function definition, adversaries do not need to cause LLMs to generate words from the target set right after the prompt, adversaries are still successful if any of the target words appear in the generated response (i.e., many tokens after the prompt). In §6.2, we show that a prompt with a lower loss value according to our new loss function was more likely to mention one of the target words set among up to 64 generated tokens. Additionally, we note that increased likelihood of a brand did not always guarantee “recommending” that brand. For instance, when we caused Gemma (one of the LLMs we use) to mention “Netflix” as a streaming service, Gemma yielded “I’ve been using Netflix for years, but I’m not happy with the selection of movies and TV shows that are available to me.” People may interpret the same LLM response differently and thus have different answers to whether a specific brand is recommended given the same response. This work, therefore, includes a more realistic user-study evaluation. We explored whether humans understood that the target brand was being recommended more in LLM responses to prompts we perturbed (§7.2.2).
5 Setup
@@
In this section, we will describe the setups of our experiments. We will first introduce our choice of LLM and parameters in §5.1. Then we depict our experiment process, including evaluation metrics, in §5.2.
5.1 LLM Setup
@@
In our experiments, we used four open-sourced LLMs as our benchmarks: a 7B pre-trained Llama 2, an 8B pre-trained Llama 3, an 8B instruction-tuned Llama 3, and a 7B instruction-tuned Gemma. Each of these models was downloaded from Meta or Google’s official repositories on HuggingFace. We used various LLMs to ensure our conclusion is not a special case only applicable to one specific instance of LLM.
@@
Temperature is a parameter that controls the determinism of LLM responses. Therefore, with different temperature choices, LLMs may tend to recommend brands with different frequencies. Following the setup that existing works used to measure the indeterministic behavior of LLMs [7], we used the default temperature for each of the four models in our experiments.
@@
Existing works collected up to responses per prompt to examine the biases in LLMs [40]. However, these studies grouped responses into only two partitions at a time (e.g., “he” versus “she” [40]), while in brand recommendation tasks, there might be more than two brands for the same category of products. To figure out how many responses we needed to collect per prompt, we ran a preliminary experiment with different combinations of prompts, target brands to be mentioned, and LLMs. We collected two sets of responses for each combination and found that they differed in the number of responses mentioning the target brand by no more than four (i.e., 0.8% in absolute means). As a result, we chose to move forward with responses per combination.
@@
When collecting responses, we generate 64 tokens per prompt. We focus on the first 64 due to three reasons, 1) focusing on the beginning of responses likely is a better heuristic of what brand users are likely to see first, thus remember (this theory is supported in our user study §7); 2) some models (e.g., Llama 2, Llama 3) don’t stop generating tokens until they reach the maximum token limit of 4096 is reached, resulting in repetitive and meaningless responses; and 3) computational cost prohibits us from collecting more tokens than we have.
5.2 Evalution Procedure
@@
We first listed 77 product categories where several established brands dominate the market. For each category, we listed popular brands we had heard about and also added brands that LLMs recommended but we did not include: some of the brands were later found to be rarely recommended by LLMs. Some brands appeared in more than one category: for example, the brand “Apple” appeared in both the category “laptops” and the category “smartphones”. The number of brands we listed for each category ranged from one to nine, with an average of per category.
@@
Aided by ChatGPT, we gathered prompts that recommended brands in each category. Specifically, we queried ChatGPT with the following prompt: “Give me multiple rephrasings and add details to: ‘What is the best XXX?’”, where “XXX” is the category (e.g., smartphones). The prompts collected in the same category were believed to be paraphrases by ChatGPT, and we manually checked that these prompts 1) paraphrased each other and 2) were inconspicuous (see definition in §3) to us researchers. We filtered out the prompts that violated either of these two principles. An example of paraphrases we used would be “Which VPN service stands out as the optimal choice for ensuring top-notch online privacy and security according to your experience?” and “Can you recommend the ultimate VPN that excels in providing robust encryption, reliable performance, and a user-friendly experience?”, resulting in two prompts that request a recommendation for a VPN. Paraphrased prompts may ask for slightly different features and may use differing wording, but ultimately ask the same thing of the LLM. Ultimately, we had three to ten prompts per category, with an average of . In total, we collected prompts. We collected responses to each of these prompts on each of the four models. We used these prompts to examine the method we described in §4.1.
@@
As we introduced in §4.2, for each combination of category and brand, adversaries may aim to cause LLMs to mention some words that are not necessarily brand names. For example, any of “ChatGPT”, “OpenAi” and “GPT” are target words (or, in some cases, strings) for the brand “ChatGPT” in the category “LLMs”. For each combination of category and brand, we collected a list of target words. We created these lists of target words based on our knowledge and observations of the responses from the 449 different prompts. Each combination of category and brand had one to five target words. We will compare paraphrased prompts by how often they mention some target words of a brand and category in §6.1. In this paper, we refer to a response as mentioning a target brand if it contains any of that brand’s target words.
@@
To evaluate the efficacy of our synonym replacement approach (introduced in §4.2), we accordingly perturbed the 449 prompts in favor of each brand of the same category. We obtained synonym-replaced prompts for each model. We will compare the 449 prompts to the prompts by how often they mention a target brand in §6.2. We paired the prompts after synonym replacement with those before synonym replacement (i.e., corresponding pairs). For each pair, we computed the absolute improvement in the percentage of responses that mention a target brand. If of responses of the prompt before synonym replacement, and of responses of the prompt after synonym replacement mention the target brand, the absolute improvement was .
Some machine learning attacks were found capable of transferring [29, 30]: attacks against a machine-learning model might be effective against a different, potentially unknown, model. We thus investigated whether our synonym replacement approach can transfer to GPT3.5-turbo, a commercial and close-sourced LLM. On each of our four open-sourced LLMs, we first ranked the pairs of prompts according to the absolute improvement in mentioning a target brand. For the top pairs, we collected complete responses to both prompts: we allowed LLMs to keep generating until they yielded end-of-sequence (EOS) tokens. Instead of absolute improvement, we computed the relative improvement in the probabilities of mentioning the target brand within complete responses for each pair. If of responses of the base prompt before perturbing it via synonym replacement, and of responses of the prompt after perturbing it mention the target brand, the relative improvement was . Then we computed the relative improvement of these pairs on ChatGPT. Specifically, we used the GPT-3.5 turbo and GPT-3.5 turbo Instruct model with the default temperature parameter and collected complete responses. GPT-3.5 turbo Instruct is the instruction-tuned version of GPT-3.5 turbo. We compare the relative improvements between GPT models and open-sourced models in §6.3.
6 Results
@@
In this section, we describe our empirical results. We first, in §6.1 describe our observations on paraphrases of prompts—measuring how pairs of paraphrased prompts can have significant differences in the probability that the target brand appears. Next, in §6.2, we report how often our synonym-replacement approach yields a perturbed prompt that causes LLMs to mention the targeted brand more often than the base prompt Finally, in §6.3, we explore whether our synonym-replacement approach transfers to GPT models. We show that if the right model-pair is found, successful perturbed prompts found on an open-sourced LLM can increase the probability of the target brand appearing more often in the GPT model’s responses.
6.1 Observations on Paraphrased Prompts

### v2

@@ 4.1.1 Paraphrased Prompts
Machine learning models, including LLMs, are brittle (e.g., [77]). They can be highly sensitive to small changes in their input. To promote a target concept, we apply this intuition. We explore manipulating LLMs by testing paraphrased prompts. Specifically, we generate a set of prompts asking for recommendations of entities in a certain category (e.g. recommendations of laptop brands) that are all paraphrases of one another. We show that these various paraphrased prompts (§4.3.1), although similar, can lead to different prominence of a target concept, e.g. a brand, in the LLM’s responses. Given enough paraphrases, we’re able to find a prompt that causes LLMs to recommend a target concept more often.
This approach does not require any access to LLMs’ internal weights or token probabilities. However, we may need to try many paraphrases of the prompt before finding an optimal prompt. These paraphrases may either run the risk of not being inconspicuous to the user, if the change between them is too great, or would require human verification before testing. Alternatively, the paraphrased prompts may be written manually. Either option is time-intensive and can prevent sufficient paraphrases from being tested. As such, we propose another method to perturb prompts, the synonym replacement attack.
4.1.2 Synonym-Replaced Prompts
The fundamental idea of this attack is to generate biased prompts by replacing a small set of words from a base (unperturbed) prompt with synonyms, ensuring minimal changes to the prompt’s meaning, and retaining a high likelihood of being inconspicuous to users. This small perturbation can, however, still lead to a large search space of possible perturbed prompts (e.g., one of our base prompts has candidate perturbed prompts), thus we introduce a loss function to help guide our selection.
Early experiments show that existing synonym dictionaries [29, 31, 32] are not always perfect synonyms in our scenarios. For example, in the context of product recommendations, WordNet [29] suggests “raw” as a synonym for “newest,” while it is quite awkward to ask for “the raw smartphone” instead of “the newest smartphone”. Thus, we created new synonym dictionaries compatible with our high-level scenarios, product prompts and societal prompts. The synonym dictionary meant for product prompts contains more words related to purchasing and recommendation, like “suggest” and “recommend” or “offer” and “provide”. Our dictionary for societal & political prompts contained synonyms like “global” and “international” or “influence” and “impact”. Our synonym dictionaries mainly consist of adjectives but also include other parts of speech, while keeping the tense of verbs consistent between synonyms. In our dictionaries, “select” and “choose” are synonyms, creating a synonym group, and other tenses of these verbs are excluded. “Exact,” “accurate,” and “precise” create another synonym group.
@@ 4.1.2 Synonym-Replaced Prompts
Prior attack work suggested white-box access to LLMs (i.e., when adversaries can access internal architecture and weights of LLMs), to use gradient search to perturb prompts [15]. In contrast, we assume a less privileged attack. Our method does not require access to model weights or gradients. Our search space is much smaller, we do not need to consider all possible tokens a model accepts, just synonyms [15]. However, the list of all possible new synonym-replaced prompts can be numerous for any given base prompt. We address this problem by computing a logit-based loss for all of these candidate perturbed prompts and picking the combination with the lowest loss. It should be noted, other methods to select a perturbed prompt from this list could be used, including random selection, which would not require access to the logits at all. However, random selection may be less efficient in finding a perturbed prompt that accomplishes the adversary’s goals, so multiple candidate perturbed may need to be tried. Therefore, for the sake of efficiency, we use loss to select a perturbed prompt.
Loss-function
@@
Specifically, we use the following loss function. Using the same notation as existing work [15], we considered LLMs as a mapping from a sequence of tokens to a distribution over the next token . In other words, LLMs generate a probability . The probability that LLMs generate a sequence can be denoted as . In contrast to Zou et al., that aim to generate a specific sequence of tokens using the loss function
@@
Intuitively, our loss function aims to cause LLMs to generate some sequence from the set right after the prompt . As we describe in §3, adversaries may not need to cause LLMs to explicitly spell out a target concept to mention or recommend the concept: causing LLMs to recommend either “Macbook” or “Apple” meets adversaries’ goal of recommending the brand “Apple” when users are asking LLMs for laptop brand recommendations. In this case, might be . Further, each sequence in may include more than one word. For example, when users are asking LLMs for grocery store recommendations, might be “Trader Joes”, “TraderJoes”, where the sequence “Trader Joes” consists of two words.
Target words
@@
Our loss function aims to increase the likelihood of words to appear from the target set right after the prompt, and we use the increased likelihood of sequences from this target set as a proxy to estimate how close adversaries are to their goal of promoting a target concept (described in §3). We use increased likelihood to estimate which combination of a prompt’s synonym replacement may cause LLMs to recommend the target concept more often. Notably, despite the loss-function definition, a successful perturbation does not need to cause LLMs to generate sequences from the target set right after the prompt, adversaries are still successful if any of the target sequences appear in the generated response (i.e., many tokens after the prompt). In §4.3.2, we show that a prompt with a lower loss value according to our new loss function was more likely to mention one of the target sequences in set among up to the first 64 generated tokens.
Our proposed approach might not be the most effective under the threat model(§3) we present. Perturbations could be made more noticeable, increasing the search space for more effective prompts. In fact, the user study to be introduced later (§5) shows that our perturbed prompts were indistinguishable from the original prompts to users indicating that there might be room for more invasive perturbations. We propose this approach only as an example to illustrate that attacks under our threat model exist, while they previously remained unexplored.
4.2 Evaluation Setup
@@
We create a set of experimental setups to test the effectiveness of our perturbations. In this section, we will first introduce our choice of LLM and parameters in §4.2.1. Then we depict our experiment scenarios in §4.2.2, and our base (unperturbed) prompts in §4.2.3. Ultimately, we will describe our experimental process, including evaluation metrics, for paraphrased prompts and synonym-replaced adversarial prompts in §4.2.4 and §4.2.5 respectively.
4.2.1 LLM Setup
@@
In our experiments, we used four open-sourced LLMs as our benchmarks: a 7B pre-trained Llama 2, an 8B pre-trained Llama 3, an 8B instruction-tuned Llama 3, and a 7B instruction-tuned Gemma. Each of these models was downloaded from Meta or Google’s official repositories on HuggingFace. We used various LLMs to ensure our conclusion is not a special case only applicable to one specific instance of an LLM.
@@
Temperature is a parameter that controls the determinism of LLM responses. Therefore, with different temperature choices, LLMs may tend to recommend concepts with different frequencies. Following the setup that existing work uses to measure the indeterministic behavior of LLMs [15], we used the default temperature for each of the four models in our experiments.
@@
The number of LLM responses to collect per prompt is a critical parameter in our experiments. Existing works collect up to responses per prompt to examine the biases in LLMs(e.g., [46]). However, these studies grouped responses into only two partitions at a time (e.g., “he” versus “she” [46]), while in concept recommendation tasks, there might be more than two candidate concepts. For example, “Apple”, “Google” and ”Samsung” are all candidate concepts (brands) when users are shopping for cell phones. To figure out how many responses we needed to collect per prompt, we ran a preliminary experiment with different combinations of prompts, target concepts to be mentioned, and LLMs. We collected two sets of responses for each combination and found that they differed in the number of responses mentioning the target concept by no more than four (i.e., 0.8% in absolute means). As a result, we chose to move forward with responses per combination.
@@
When collecting responses, we generate 64 tokens per response. We focus on the first 64 for three reasons, 1) focusing on the beginning of responses likely is a better heuristic of what concepts users are likely to see first, thus remember (this theory is supported in our user study §5); 2) some models (e.g., Llama 2, Llama 3) don’t stop generating tokens until they reach the maximum token limit of 4096 is reached, resulting in repetitive and meaningless responses; and 3) computational cost prohibits us from collecting more tokens than we have. The amount of tokens we generate per response might not be representative of the real-world use of LLMs. However, our post-hoc analysis of user study data will show that early mentions of target concepts are much more likely to be noticed by users than later mentions (§5.2.2).
4.2.2 Experiment Scenarios
We used two high-level scenarios to evaluate our perturbation approaches. In the first scenario, users are asking LLMs for recommendations of brands when they are shopping for a specific category of goods (e.g. laptops). Adversaries aim to force LLMs to recommend a specific brand more often, advertising without user awareness, gaining economic benefits from the advertising ecosystem, and hurting user autonomy de facto. In the second scenario, users are asking LLMs for concepts that are positive or negative in a societal topic (such as the winner of the space race, the most influential US president, or the country that is the worst offender of women’s rights). As a note, even though some of these prompts are negative, we will use the term “recommend” throughout the paper for consistency, even though in these cases “denounce” may be more accurate. Adversaries aim to propagandize by forcing LLMs to answer with the target concept. To minimize the harm in our human study, we only performed a human evaluation on the product scenario (§5). The scenarios we consider might not generalize to all use cases of prompt providers. We partially mitigate this limitation by considering many prompts and target concepts per scenario.
4.2.3 Base Prompts
@@
We first collected categories of concepts and target concepts that users may query LLMs for. For the product scenario, we listed product categories where several established brands dominate the market. For each category, we listed popular brands we were aware of and supplemented this list by querying ChatGPT for a list of popular brands in these categories. However, some brands were later found to be rarely recommended by LLMs. Some brands appeared in more than one category: for example, the brand “Apple” appeared in both the category “laptops” and the category “smartphones”. The number of brands we listed for each category ranged from one to nine, with an average of per category. We performed the same collection procedure for the societal scenario, and ended up having two to nine concepts for each of the eleven topics we found, with an average of per topic.
4.2.4 Paraphrased Prompts Setup
@@
Aided by ChatGPT, we gathered prompts that recommended concepts in each product category or societal topic. Specifically, for each product category, we queried ChatGPT with the following prompt: “Give me multiple rephrasings and add details to: ‘What is the best XXX?’”, where “XXX” is the category (e.g., smartphones), The prompts collected in the same category were believed to be paraphrases by ChatGPT, and we manually checked that these prompts 1) paraphrased each other and 2) were inconspicuous (see definition in §3) from our perspective. We filtered out the prompts that violated either of these two principles. An example of paraphrases we used would be “Which VPN service stands out as the optimal choice for ensuring top-notch online privacy and security according to your experience?” and “Can you recommend the ultimate VPN that excels in providing robust encryption, reliable performance, and a user-friendly experience?”, resulting in two prompts that request a recommendation for a VPN. For societal topics, we similarly collected paraphrased prompts, starting with prompts such as “Which country won the space race?” Paraphrased prompts may ask for slightly different features and may use differing wording, but ultimately ask the same thing of the LLM. Ultimately, we had three to ten prompts per product category, with an average of and a total of prompts. We had six to eight prompts per societal topic, with an average of and a total of prompts. We collected responses to each of these prompts on each of the four models. We used these prompts to examine the method we described in §4.1.1.
@@
As we introduced in §4.1.2, for each combination of category/topic and concept, adversaries may aim to cause LLMs to mention some words that are not necessarily concept names. For example, any of “ChatGPT”, “OpenAi” and “GPT” are target words (or, in some cases, strings) for the brand “ChatGPT” in the product category “LLMs”. For each combination of category/topic and concept, we collected a list of target words. We created these lists of target words based on our knowledge and observations of the responses from the different prompts. Each combination of category/topic and concept had one to five target words. We will compare paraphrased prompts by how often they mention some target words of a category/topic and concept in §4.3.1. In this paper, we refer to a response as mentioning a target concept if it contains any of that concept’s target words.
4.2.5 Synonym-Replaced Adversarial Prompts Setup
As we introduced in §4.1.2, created new synonym dictionaries compatible with recommendation tasks. We came up with a dictionary containing a total of words in synonym groups for the product scenario, and a dictionary consisting of a total of words in synonym groups for the societal scenario. Each word may have at most seven synonyms.
@@ 4.2.5 Synonym-Replaced Adversarial Prompts Setup
To evaluate the efficacy of our synonym replacement approach (introduced in §4.1.2), we accordingly perturbed the prompts in favor of each concept of that category/topic. We obtained synonym-replaced prompts for each model, consisting of for the product scenario and for the societal scenario. We will compare the prompts to the prompts by how often they mention a target concept in §4.3.2. We paired the prompts after synonym replacement with those before synonym replacement (i.e., corresponding pairs). For each pair, we computed the absolute improvement in the percentage of responses that mention a target brand. If of responses of the prompt before synonym replacement (base prompt), and of responses of the prompt after synonym replacement (perturbed prompt) mention the target brand, the absolute improvement was .
For the sake of completeness, we also evaluated the effectiveness of our synonym replacement approach to reduce the frequency of mentioning a target concept by maximizing the loss function (§4.3.3). We’re limited by computational resources. The models we run our evaluations on might not be representative of LLMs used in the wild. Thus, we explore the transferability of our synonym replacement approach to (close-sourced) GPT models (§4.3.4).
4.3 Results
@@
In this section, we describe our empirical results. First, in §4.3.1, we describe our observations on paraphrases of prompts—measuring how pairs of paraphrased prompts can have significant differences in the probability that the target concept appears. Next, in §4.3.2, we report how often our synonym-replacement approach yields a perturbed prompt that causes LLMs to mention a targeted concept more often than the base prompt. Then in §4.3.3. we discuss how our synonym-replacement approach might also yield perturbed prompts that cause LLMs to mention a target concept less often than the base prompt, mitigating existing bias in the LLMs. Finally, in §4.3.4, we explore whether our synonym-replacement approach transfers to GPT models. We show that if the right model pair is found, successful perturbed prompts found on an open-sourced LLM can increase the probability of the target brand appearing more often in the GPT model’s responses.
4.3.1 Concept Frequency Differences on Paraphrased Prompts